OpenStudy (anonymous):

PLEASE HELP EASY QUESTION!! Create your own binomial expression with a radical in the second term.

OpenStudy (john_es):

May be, \[a+\sqrt{b}\]

OpenStudy (anonymous):

with numbers

OpenStudy (john_es):

\[2+\sqrt{3}\]

OpenStudy (anonymous):

can u please give me another one than 2+radical 3 and 2+radical 5?

OpenStudy (john_es):

\[3+4\sqrt[4]{3}\]

OpenStudy (anonymous):

Create your own binomial expression with a radical in the second term. Part 1: Identify its conjugate and explain, in complete sentences, why it is the conjugate. (1 point) Part 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why?

OpenStudy (anonymous):

@John_ES this is the question

OpenStudy (anonymous):

@John_ES ??????

OpenStudy (john_es):

Binomial expression, \[4+\sqrt{5}\] Its conjutage, \[4-\sqrt{5}\] It is its conjugate because it has the radical with the opposite sign. Product, \[(4+\sqrt{5})(4-\sqrt{5})=4^2-(\sqrt{5})^2=16-5=11\]